OpenStudy (anonymous):

Solve. 7^(x - 5) = 7^(4x + 7)

sam (.sam.):

x-5=4x+7

sam (.sam.):

Solve

OpenStudy (lgbasallote):

rules of exponential thingy...when a^m = a^n...m=n so x-5 = 4x +7